# Editable Fairness: Fine-Grained Bias Mitigation in Language Models

### v1

@@ 7 Conclusion and Limitation
In this paper, we explore the fine-grained bias mitigation paradigm, which focuses on individual social biases rather than group differences. The exploration has been developed from two aspects. We first establish a novel debiasing benchmark BiaScope to finely measure the effects of debiasing, which not only evaluates the fairness on the social biases, but also the retention of individual commonsense knowledge. Furthermore, we introduce the first editable bias mitigation framework FAST, which is capable of locating and mitigating individual social biases precisely. Experiments have validated that existing group-invariant debiasing techniques severely affect the retention of commonsense knowledge, while demonstrating the superiority of FAST in both bias mitigation and knowledge maintenance. Extensive experiments across various models and datasets further demonstrate its scalability, robustness, and lightweight. Our findings offer significant implications for future debiasing research. Broader impact is discussed in Appendix F.
@@
We acknowledge the presence of certain limitations. First, in this paper, we construct our new datasets leveraging GPT-4. Although human validation is performed to ensure the reliability of the data, GPT-4 may suffer from the limitations of its internal knowledge, potentially introducing blind spots into our benchmark. Second, the memory mechanism of LLMs is still under exploration, while we assume that FFN layers are responsible for storing biased knowledge based on previous observations [25, 46, 26, 8]. Third, debiasing larger models, as shown in Table 5.2, is more challenging and will guide our future research, which constitutes our future direction. Besides, social bias in open language generation or dialogue represents another critical scenario for mitigating bias [66], which constitutes one of our future research endeavors.
@@ Appendix A BiaScope Benchmark Construction
A.1 Metrics
Stereotype Score (SS) is the most straightforward measure for the bias within the debiased model [51, 53]. It computes the percentage of knowledge for which a model assigns the biased object as opposed to the unbiased object. The evaluation of SS is conducted according to the following criteria:
where $\mathcal{G}^{*}$ is the debiased model.
Language Modeling Score (LMS), employed in StereoSet [51], has been utilized. Based on the knowledge pairs in $\Omega_{S}$, we select an irrelevant $o_{ir}$ to form $k_{ir}=(s,r,o_{ir})$. LMS represents the percentage that a model that prefers a relevant association (either the stereotypical association or the anti-stereotypical association) as opposed to an irrelevant association. The evaluation of LMS is conducted according to the following criteria:
$\textbf{LMS}({\mathcal{G}},\Omega_{S})$ $=\mathbb{E}_{(k_{1},k_{2})\in\Omega_{S}}\mathbbm{1}\{\mathcal{P}_{\mathcal{G}}[k_{1}]>\mathcal{P}_{\mathcal{G}}[k_{ir}]\}$ (9)
Ideal Context Association Test Score (ICAT) is proposed by [52] combine both LMS and SS by $\text{ICAT}=\text{LMS}*\text{min}(\text{SS},100-\text{SS})/50$. It represents the language modeling ability of a model while behaving in an unbiased manner.
A.2 Dataset Construction Details
To ensure the quality and diversity of the generated data, meticulous human validation was performed. We employed four human evaluators, all of whom possess good English proficiency, to annotate and verify the feasibility of the datasets generated by GPT-4.
@@ A.2 Dataset Construction Details
Paraphrased dataset. For each knowledge pair within $\Omega_{S}$, we paraphrase the prompts combining $(s,r)$ with the same semantic expression. We hired 2 undergraduate students, all with good English proficiency. We first asked the students to manually paraphrase ten pieces of biased knowledge into semantically similar ones. Then, the manually paraphrased samples were combined with the prompt as context for GPT-4 generation. After generation, we performed sample checks on 10% of the data for each dataset. In these samples, the agreement on successful generation reached 100%.
@@
Retention dataset. We construct $\Omega_{R}$ by collecting commonsense facts related to the sensitive attributes, such as “Jesus’ resurrection is commemorated by [Christians] when they celebrate Easter.” We initially created alternative facts by prompting the GPT-4 API. We then asked the students to manually validate every generated fact, ensuring that each fact in the retention dataset constitutes reasonable commonsense knowledge rather than bias.
A.3 Diversity and Challenge Analysis of BiaScope
@@
Appendix B Method
B.1 Preliminaries
Considering a transformer-based Large Language Model (take a decoder-based transformer as an example). The model takes into the input sequence $(x_{1},...x_{t-1})$ and predicts the probability of the next token ($x_{t}$). The updates in a transformer block can be formulated as follows:
$h_{t}^{(l)}=h_{t}^{(l-1)}$ $+a_{t}^{(l)}+m_{t}^{(l)},$ (11)
where $\text{where }a_{t}^{(l)}$ $=\text{Att}(h_{1}^{(l-1)},h_{2}^{(l-1)},...,h_{t}^{(l-1)}),$ (12)
$m_{t}^{(l)}$ $=W_{2}\sigma(W_{1}\gamma(a_{t}^{(l)}+h_{t}^{(l-1)})),$ (13)
where $m_{t}^{(l)}$ and $a_{t}^{(l)}$ represents the output of the $l^{t}h$ feed-forward network layer and self-attention layer.
B.2 Locate Biased Knowledge
In this section, we provide a complete illustration of our Step 1 in Figure 6.
Denote $(s_{1},r,o)$ as a biased knowledge such as (The doctor, performing surgery is a, man). $(s_{2},r,o)$ is the counterfactual knowledge (i.e., $s_{2}$ is The nurse). Our biased knowledge localization is performed in three steps, with a complete illustration in Figure 6 in the Appendix:
Biased run: We pass the prompt $(s_{1},r)$ into the model and collect all hidden states $\{h^{(l)}_{i}$ |||| $i\in[1,T]$, $l\in[1,L]\}$ where $T$ is number of tokens and $L$ is number of layers.
Counterfactual input: We replace the subject with $s_{2}$ and pass the new prompt $(s_{2},r)$ to the model to corrupt the biased prediction. Hidden states corresponding to the subject token(s) $\hat{i}$ will be updated with $h_{\hat{i}}^{(0)}(s_{1}\rightarrow s_{2})$.
Restoration run: Towards certain layer $\hat{l}$ in the model, we hook the biased states $h_{\hat{i}}^{(\hat{l})}$ at subject token(s) $\hat{i}$ and perform the counterfactual run. Then we calculate the recovery degree of biased prediction, which indicates the causal effect of $\hat{l}$ to biased prediction. The layer with highest causal effect will be selected as the decisive layer.
Causal effect. Denote $\mathcal{P}[o]$, $\mathcal{P}^{*}[o]$ as the probability of biased prediction and counterfactual prediction. Let $\mathcal{P}^{*}(h_{\hat{i}}^{(\hat{l})})[o]$ denotes the probability of counterfactual prediction with restoration of the biased states $h_{\hat{i}}^{(\hat{l})}$. The indirect causal effect (IE) of a certain layer can be calculated by $\text{IE}=\mathcal{P}^{*}(h_{\hat{i}}^{(\hat{l})})[o]-\mathcal{P}^{*}[o]$.
Figure 6: Illustration of our debiasing framework.
@@ C.1 Experiment details
Dataset. We collect biased knowledge related to three domains (gender, race, and religion) from six existing datasets (StereoSet [51], Crows-Pairs [53], WEAT [6], WinoBias [75], Winogender [59] and BEC-Pro [2]). These datasets have been benchmarked to detect biases within Language Models (LLMs). The statistics of our constructed knowledge base can be referred to Table 11, with a detailed description referred to in the following.
StereoSet [51] employs a methodology to evaluate a language model’s propensity for stereotypical associations. The procedure is essentially a fill-in-the-blank challenge, where the model is given a sentence with a missing word and must select from a stereotypical word, an anti-stereotypical word, or an irrelevant word.
CrowS-Pairs [53] constitutes a dataset featuring intrasentential minimal pairs. Each pair comprises one sentence depicting a socially disadvantaged group in a manner that either conforms to or contradicts a stereotype, and another sentence that is slightly altered to reference a contrasting, advantaged group. The language model’s task involves assessing the probability of masked tokens that are exclusive to each sentence within these pairs.
WEAT [6] is comprised of word sets that pertain to either attributes or targets. It evaluates the associations between concepts of social groups (for instance, masculine and feminine terms) and neutral attributes (such as terms related to family and occupation).
Winogender [59] and Winobias [76] are designed to assess gender-based stereotypical associations with various occupations. In some instances, these evaluations involve associating gender-specific pronouns with occupations that are stereotypically linked to that gender. In other cases, the task is to associate pronouns with occupations that are typically considered non-stereotypical for that gender.
BEC-Pro (The Bias Evaluation Corpus with Professions) [2] is a tool for assessing gender biases in the context of occupations. It comprises 5,400 sentences, each generated from a template that includes a term denoting a person and one of 60 professional terms. During the evaluation process, both the person-related and professional words in these sentences are masked for analysis.
@@
Baselines. We consider the following debiasing techniques as baselines. The techniques can be grouped into two categories. (1) Fine-tuning: Counterfactual Data Augmentation (CDA) [^1] We use the reproduction of CDA, Dropout, SentenceDebias, INLP and Self-Debias provided by https://github.com/McGill-NLP/bias-bench [78] involves re-balancing a corpus by swapping bias attribute words (e.g., he/she) in a dataset. The re-balanced corpus is then often used for further training to debias a model. Dropout [70] proposes to increase the dropout parameters and perform an additional phase of pre-training to debias. SentenceDebias [43] proposes to obtain debiased representation by subtracting biased projection on the estimated bias subspace from the original sentence representation. Iterative Nullspace Projection (INLP) [58] is also a projection-based debiasing technique to remove protected property from the representations. MABEL [^2] We use the debiased models provided in https://github.com/princeton-nlp/MABEL/ [33] mitigates Gender Bias using Entailment Labels. (2) Prompt-tuning: Auto-debias [^3] We use the debiased models provided in https://github.com/Irenehere/Auto-Debias [27] proposes to directly probe the biases encoded in pre-trained models through prompts, then mitigate biases via distribution alignment loss. (3) Post-hoc: Self-Debias [61] proposes to leverage a model’s internal knowledge to discourage it from generating biased text. FMD [7] proposes a machine unlearning-based strategy to efficiently remove the bias in a trained model. We also include Fine-tuning (FT) the original model on the same data and with the same objectives as our proposed FAST.
@@
Implementation details. Bias mitigation is conducted over the collected biased knowledge in Section 4.1. We utilize two-layer fully connected neural networks with the ReLU activation function as the fairness stamp. The hidden dimension is 1024. The batch size is set to 4. We use Adam optimizer with a learning rate of 0.1. We train each batch for 20 iterations. $\alpha$ is set to be 40 and $\beta$ is 0.1. The model is trained on 8 RTX 3090 GPUs with 24G memory. We utilize pre-trained backbone models in the Huggingface Transformers library [71].
Source/domain gender race religion
StereoSet 771 2976 247
Crows-Pairs 262 516 105
Winogender 60 - -
BEC-Pro 5400 - -
@@
C.2 Knowledge Locating Results

### v2

@@ 6 Conclusion
In this paper, we explore the fine-grained bias mitigation paradigm, which focuses on individual social biases rather than group differences. The exploration has been developed from two aspects. We have developed a new debiasing benchmark, BiaScope, which evaluates not only fairness regarding social biases but also the preservation of individual commonsense knowledge. Furthermore, we introduce the first editable bias mitigation framework FAST, which is capable of locating and mitigating individual social biases precisely. Experiments have demonstrated the superiority of FAST in both bias mitigation and knowledge maintenance. Extensive experiments across various models and datasets further demonstrate its scalability, robustness, and lightweight. Our findings offer significant implications for future debiasing research.
Acknowledgement
This work is supported by the National Natural Science Foundation of China (Grant No. 62476241), the Natural Science Foundation of Zhejiang Province, China (Grant No. LZ23F020008), and the Zhejiang University-Angelalign Inc. R&D Center for Intelligent Healthcare.
Limitation
@@ Acknowledgement
We acknowledge the presence of certain limitations. First, in this paper, we construct our new datasets leveraging GPT-4. Although human validation is performed to ensure the reliability of the data, GPT-4 may suffer from the limitations of its internal knowledge, potentially introducing blind spots into our benchmark. Second, the memory mechanism of language models is still under exploration, while we assume that FFN layers are responsible for storing biased knowledge based on previous observations Geva et al. (2020); Meng et al. (2022a); Geva et al. (2022). Third, debiasing larger models, as shown in Table 4, is more challenging and will guide our future research, which constitutes our future direction. Besides, social bias in open language generation or dialogue represents another critical scenario for mitigating bias Wan et al. (2023), which constitutes one of our future research endeavors.
Potential Risks
With the widespread application of language models, the emphasis on fairness has significantly increased, requiring language models to treat individuals from different backgrounds fairly. However, language models trained on large datasets inevitably exhibit certain biases during the pre-training phase. In this paper, we propose a promising solution that mitigates unfairness in language models while not compromising capability, which is of great significance for deploying fair and reliable language models. This research utilizes publicly available datasets and performs human validation on the created datasets, ensuring that all data complies with privacy regulations and has been anonymized where necessary. Our aim is to promote the responsible and fair use of LLMs to enhance accessibility and automation, while advocating for ethical AI development. Our study does not involve human subjects or violate legal compliance. At present, no additional potential risks have been identified.
@@ Appendix A BiaScope Benchmark Construction
A.1 Datasets
We collect biased knowledge related to three domains (gender, race, and religion) from six existing datasets (StereoSet Nadeem et al. (2020b), Crows-Pairs Nangia et al. (2020), WEAT Caliskan et al. (2017), WinoBias Zhao et al. (2018), Winogender Rudinger et al. (2018) and BEC-Pro Bartl et al. (2020)). These datasets have been benchmarked to detect biases within Language Models. The statistics of our constructed knowledge base can be referred to Table 7, with a detailed description referred to in the following.
StereoSet Nadeem et al. (2020b) employs a methodology to evaluate a language model’s propensity for stereotypical associations. The procedure is essentially a fill-in-the-blank challenge, where the model is given a sentence with a missing word and must select from a stereotypical word, an anti-stereotypical word, or an irrelevant word.
CrowS-Pairs Nangia et al. (2020) constitutes a dataset featuring intrasentential minimal pairs. Each pair comprises one sentence depicting a socially disadvantaged group in a manner that either conforms to or contradicts a stereotype, and another sentence that is slightly altered to reference a contrasting, advantaged group. The language model’s task involves assessing the probability of masked tokens that are exclusive to each sentence within these pairs.
WEAT Caliskan et al. (2017) is comprised of word sets that pertain to either attributes or targets. It evaluates the associations between concepts of social groups (for instance, masculine and feminine terms) and neutral attributes (such as terms related to family and occupation).
Winogender Rudinger et al. (2018) and Winobias Zhao et al. (2019) are designed to assess gender-based stereotypical associations with various occupations. In some instances, these evaluations involve associating gender-specific pronouns with occupations that are stereotypically linked to that gender. In other cases, the task is to associate pronouns with occupations that are typically considered non-stereotypical for that gender.
BEC-Pro (The Bias Evaluation Corpus with Professions) Bartl et al. (2020) is a tool for assessing gender biases in the context of occupations. It comprises 5,400 sentences, each generated from a template that includes a term denoting a person and one of 60 professional terms. During the evaluation process, both the person-related and professional words in these sentences are masked for analysis.
Source/domain gender race religion
StereoSet 771 2976 247
Crows-Pairs 262 516 105
Winogender 60 - -
BEC-Pro 5400 - -
A.2 Instructions to GPT-4
A.2.1 Commonsense dataset
  • System
    You are a helpful assistant. Please generate text that provides specific commonsense knowledge related to the given social group, ensuring that the content is free of stereotyped biases. Produce a concise sentence that offers respectful commonsense knowledge about the group.
  • User
    Original Social Bias
    Mom
  • Assistant
    Commonsense Social Bias
    Mom generally gives birth to the child.
A.2.2 Paraphrased dataset
  • System
    You are a helpful assistant. Please paraphrase the following text related to social bias. It is important to preserve the original meaning and structure of the input as closely as possible. Please deliver a paraphrased version of the text that retains the factual content and logical sequence of the original text. Do not introduce new information or examples that are not present in the original text. Keep the paraphrase approximately the same length as the original text.
  • User
    Original Social Bias
    Mom is more likely to take care of the child.
  • Assistant
    Paraphrased Social Bias
    Mom is more likely to raise the child.
A.3 Human Validation
To ensure the quality and diversity of the generated data, meticulous human validation was performed. Five undergraduate students were recruited as human annotators, each demonstrating superior proficiency in English and adhering to stringent ethical standards. These individuals were strategically selected to embody a broad spectrum of demographic characteristics, including diverse ages, ethnic backgrounds, and cultural perspectives, with statistics detailed in Table 8. They have partaken in numerous extensive discussions with our research team to thoroughly comprehend the essential requirements for the evaluation procedure. In compliance with applicable local labor laws and regulations, these individuals are remunerated based on the number of hours worked, thus ensuring fair compensation for their contributions.
Paraphrased dataset.
@@ A.3 Human Validation
For each knowledge pair within $\Omega_{S}$, we paraphrase the prompts combining $(s,r)$ with the same semantic expression. We first asked the students to manually paraphrase ten pieces of biased knowledge into semantically similar ones. Then, the manually paraphrased samples were combined with the prompt as context for GPT-4 generation. After generation, we performed sample checks on 10% of the data for each dataset. In these samples, the agreement on successful generation reached 100%.
Annotator A B C D E
Gender Male Male Female Male Female
Age Group 18-30 18-30 18-30 30-50 30-50
Race Asian Asian Asian European South American
Religion Non-religious Buddhist Non-religious Christian Non-religious
Table 8: Annotator Information
Retention dataset.
@@
We construct $\Omega_{R}$ by collecting commonsense facts related to the sensitive attributes, such as “Jesus’ resurrection is commemorated by [Christians] when they celebrate Easter.” We initially created alternative facts by prompting the GPT-4 API. We then asked the students to manually validate every generated fact, ensuring that each fact in the retention dataset constitutes reasonable commonsense knowledge rather than bias.
A.4 Instructions to Human Annotator
A.4.1 Commonsense dataset
[Task] You are required to conduct a detailed inspection of the generated commonsense dataset. Please pay special attention to the following points:
1. Confirm whether each fact conforms to generally accepted commonsense.
2. Ensure that the fact descriptions do not contain any form of bias or discriminatory language. [Procedures] Carefully read each commonsense fact. Evaluate the facts based on their commonsense validity, and absence of bias. Label each piece of data with tags such as ‘Pass’, ‘Fail’, or ‘Uncertain.’
[Notes] If you have any doubts about the assessment of a fact during the annotation process, please contact us.
A.4.2 Paraphrased dataset
[Task] You need to conduct a detailed inspection of the generated paraphrased dataset. Please pay special attention to the following aspects:
1. Verify whether the paraphrased content maintains the core semantics of the original knowledge pairs.
2. Check whether the rewritten texts are fluent and natural to read, with no grammatical errors or unnatural expressions.
[Procedures] Read each paraphrased piece of knowledge. Compare with the original data to assess whether the rewritten content meets the requirements specified in Task. Label each data item as ‘Pass’, ‘Fail’, or ‘Uncertain’.
[Notes] If you have any doubts about the assessment of a fact during the annotation process, please contact us.
A.5 Diversity and Challenge Analysis
@@
Appendix B Related Works
Bias Mitigation in Pre-trained Language Models.
The increasing deployment of chatbots driven by large language models (LLMs) has heightened concerns over fairness. Issues related to fairness in LLMs can have dire outcomes, such as the amplification of bias, discrimination, and detrimental effects on marginalized groups. Consequently, substantial efforts are being made to assess and address biases in large language models Gallegos et al. (2023); Li et al. (2023b); Fan et al. (2024a); Luo et al. (2024). Several approaches have been proposed for debiasing pre-trained language models, which can be grouped into two categories: (1) Fine-tuning. This branch includes additional pre-training on re-balanced corpora Zmigrod et al. (2019); Webster et al. (2020) or with a contrastive objective He et al. (2022); Cheng et al. (2021), projection-based methods Liang et al. (2020); Ravfogel et al. (2020); Kaneko and Bollegala (2021); Dev et al. (2020) in the embedding space, in-training methods Han et al. (2021); He et al. (2022) and parameter-efficient fine-tuning Lauscher et al. (2021); Xie and Lukasiewicz (2023) methods. (2) Prompt-tuning. Prompt-tuning Guo et al. (2022); Yang et al. (2023); Li et al. (2023c); Dong et al. (2023) involve generating either discrete prompts or continuous prompts to mitigate social biases. There are also post-hoc approaches Schick et al. (2021); Chen et al. (2023, 2024a); Fan et al. (2024b); Chen et al. (2024b) that are deployed after the training phase to achieve effective debiasing. However, existing techniques treat social groups as interchangeable Gallegos et al. (2023) and neutralize different social groups in model inputs or outputs, while ignoring or concealing distinct facts of different social groups Hanna et al. (2020). In contrast, our method mitigates biases based on fine-grained individual biases, avoiding compromising other knowledge.
[…]
BERT 60.28 59.17 100.0 84.17
$\checkmark$ - - 47.92 49.27 52.38 66.72
$\checkmark$ $\checkmark$ - 51.76 52.06 92.86 86.93
$\checkmark$ $\checkmark$ $\checkmark$ 51.16 50.80 95.83 86.30
Knowledge Locating.
Knowledge Locating aims to interpret how knowledge is encapsulated within specific model components, including neurons, layers, or subnetworks Elhage et al. (2021); Rogers et al. (2021); Schneider and Vlachos (2021); Zeiler and Fergus (2014); Wang et al. (2022); Bolukbasi et al. (2021). Geva et al. (2020) proposes that it is the FFN layers that serve as repositories of factual knowledge, while other works Elhage et al. (2021); Hao et al. (2021) illustrate that the self-attention mechanism is instrumental in replicating information. More recent works Meng et al. (2022a); Geva et al. (2022, 2023) posit that the feed-forward components of transformer-based PLMs function akin to key-value memory systems, archiving data pertinent to specific subjects. Inspired by these works, we are the first to define social biases as a knowledge triplet (subject, relation, object), stemming from our observation that a social bias typically consists of a biased description (i.e., object) directed towards a certain social group (i.e., subject). Furthermore, we propose using a counterfactual knowledge pair to trace states by altering the social group or biased description, due to the fact that social biases represent inequitable attitudes or perceptions between social groups (e.g., male, female) regarding abilities (e.g., good at math/art).
Knowledge Editing.
Knowledge or Model Editing Sinitsin et al. (2020); De Cao et al. (2021); Dai et al. (2021) has been proposed to facilitate data-efficient modifications to model behavior while ensuring no detrimental impact on performance across other inputs. These approaches manipulate the model’s output for specific cases either by integrating external models with the original, unchanged model Mitchell et al. (2022); Murty et al. (2022); Dong et al. (2022); Hartvigsen et al. (2022); Huang et al. (2023); Zheng et al. (2023), or by altering the model parameters responsible for undesirable output Mitchell et al. (2021); Gupta et al. (2023); Hase et al. (2021); Meng et al. (2022a). The most relevant line of work is locate and edit Meng et al. (2022a, b); Dai et al. (2021); Li et al. (2023a), which suggests identifying neurons crucial to a model’s factual predictions Vig et al. (2020); Finlayson et al. (2021) and subsequently updating the feed-forward weights to edit the output. Inspired by these works, we propose the first fine-grained bias mitigation framework, which enables the nuanced calibration of individual social biases at minimal cost. This addresses the issue in existing methods where the excessive pursuit of equality leads to incorrect predictions.
@@ C.1 Experiment details
Baselines.
@@
We consider the following debiasing techniques as baselines. The techniques can be grouped into two categories. (1) Fine-tuning: Counterfactual Data Augmentation (CDA) [^3] We use the reproduction of CDA, Dropout, SentenceDebias, INLP and Self-Debias provided by https://github.com/McGill-NLP/bias-bench Zmigrod et al. (2019) involves re-balancing a corpus by swapping bias attribute words (e.g., he/she) in a dataset. The re-balanced corpus is then often used for further training to debias a model. Dropout Webster et al. (2020) proposes to increase the dropout parameters and perform an additional phase of pre-training to debias. SentenceDebias Liang et al. (2020) proposes to obtain debiased representation by subtracting biased projection on the estimated bias subspace from the original sentence representation. Iterative Nullspace Projection (INLP) Ravfogel et al. (2020) is also a projection-based debiasing technique to remove protected property from the representations. MABEL [^4] We use the debiased models provided in https://github.com/princeton-nlp/MABEL/ He et al. (2022) mitigates Gender Bias using Entailment Labels. (2) Prompt-tuning: Auto-debias [^5] We use the debiased models provided in https://github.com/Irenehere/Auto-Debias Guo et al. (2022) proposes to directly probe the biases encoded in pre-trained models through prompts, then mitigate biases via distribution alignment loss. (3) Post-hoc: Self-Debias Schick et al. (2021) proposes to leverage a model’s internal knowledge to discourage it from generating biased text. FMD Chen et al. (2023) proposes a machine unlearning-based strategy to efficiently remove the bias in a trained model. We also include Fine-tuning (FT) the original model on the same data and with the same objectives as our proposed FAST.
Implementation details.
@@
Bias mitigation is conducted over the collected biased knowledge in Section 2.2. We utilize two-layer fully connected neural networks with the ReLU activation function as the fairness stamp. The hidden dimension is 1024. The batch size is set to 4. We use Adam optimizer with a learning rate of 0.1. We train each batch for 20 iterations. $\alpha$ is set to be 40 and $\beta$ is 0.1. The model is trained on 8 RTX 3090 GPUs with 24G memory. We utilize pre-trained backbone models in the Huggingface Transformers library Wolf et al. (2020).
@@
Metrics
Stereotype Score (SS) is the most straightforward measure for the bias within the debiased model Nadeem et al. (2020b); Nangia et al. (2020). It computes the percentage of knowledge for which a model assigns the biased object as opposed to the unbiased object. The evaluation of SS is conducted according to the following criteria:
where $\mathcal{G}^{*}$ is the debiased model.
Language Modeling Score (LMS), employed in StereoSet Nadeem et al. (2020b), has been utilized. Based on the knowledge pairs in $\Omega_{S}$, we select an irrelevant $o_{ir}$ to form $k_{ir}=(s,r,o_{ir})$. LMS represents the percentage that a model that prefers a relevant association (either the stereotypical association or the anti-stereotypical association) as opposed to an irrelevant association. The evaluation of LMS is conducted according to the following criteria:
$\textbf{LMS}({\mathcal{G}},\Omega_{S})$ $=\mathbb{E}_{(k_{1},k_{2})\in\Omega_{S}}\mathbbm{1}\{\mathcal{P}_{\mathcal{G}}[k_{1}]>\mathcal{P}_{\mathcal{G}}[k_{ir}]\}$
@@
Ideal Context Association Test Score (ICAT) is proposed by Nadeem et al. (2020a) combine both LMS and SS by $\text{ICAT}=\text{LMS}*\text{min}(\text{SS},100-\text{SS})/50$. It represents the language modeling ability of a model while behaving in an unbiased manner.
C.2 Bias Knowledge Localizing Results
C.3 Qualitative Study of Bias Mitigation